# Impurity flat band states in the diamond chain

### v1

@@ Abstract
Flat band (FB) systems, i.e., models hosting dispersionless bands in their energy spectrum, have attracted great interest recently. The eigenstates of these FBs can be written as compact localized states (CLSs), meaning states having finite support only over a small, non-extensive portion of the lattice. When adjacent CLSs of a FB have spatial overlap, the subspace they span is, in general, non-orthogonal. Here, we consider precisely this scenario, in the form of a diamond chain with a finite magnetic flux per plaquette, which has a gapped midspectrum FB. After developing the general framework for projecting arbitrary operators onto a non-orthogonal subspace, we address the case of an open chain with small local impurities in the midchain plaquette. The framework is then applied to derive analytical expressions for the states of the FB affected by the impurities. The results are illustrated for three representative cases: two equal impurities, two symmetric impurities, and a single impurity. We also address the effects of diagonal disorder on the impurity states. Focusing on the equal impurities case, we show how these states effectively feel an averaged disorder determined by their spatial extension, which can be controlled by the flux. Due to this averaging effect, these states exhibit enhanced robustness to disorder. For the single-impurity case, we further reveal its exotic non-trivial topological phase, without precedent in the literature on Hermitian models, characterized by a half-integer winding number. By mapping the system onto an equivalent two-dimensional (2D) tight-binding lattice decorated with a central impurity line, the half-integer bulk invariants are shown to have a correspondence with the appearance of a single in-gap edge state when open boundary conditions (OBC) are imposed. Numerical simulations are performed to benchmark all analytical predictions, showing a near-perfect agreement for flux values away from the gap closing point. Our results lay the foundation for further inquiries on the physics of non-orthogonal bases, and are expected to be of particular relevance to the study of FB related phenomena.
I Introduction
About a decade ago, a renewed interest in FB systems [1] gave way to a new cycle of studies devoted to the topic (a more detailed chronology can be found in [2]), and it has no signs of slowing down. In this new wave of studies, the focus has gradually shifted towards the CLSs that span every FB. For one-dimensional (1D) models, it has been shown that these CLSs form an orthogonal set when the FB projector is strictly local, and a non-orthogonal one otherwise [3, 4]. Based on the form of the CLSs, FB generators were constructed for 1D [5, 6] and higher-dimensional [7, 8, 9, 10, 11] models.
@@ I Introduction
Here, we address the case of a diamond chain with a finite flux per plaquette, and introduce a perturbation in the form of weak local potentials in the midchain plaquette, as schematically illustrated in Fig. 1(a). By projecting this impurity perturbation onto the gapped zero-energy FB, only two partially overlapping CLSs have finite weight at the impurity sites. However, due to the non-orthogonal nature of adjacent CLSs, we show that the analytical treatment of the system requires one to reexpress the problem in terms of its dual basis states. This greatly simplifies the problem, by reducing it to the solutions of a decoupled dimer in a non-orthogonal dual subspace. We then derive the analytical solutions for the impurity states that are lifted from the FB as a function of the magnetic flux, focusing on the cases of two equal impurities, two symmetric impurities, and a single-impurity, where near-perfect agreement with numerical results is observed. Finally, for the single-impurity case, we further demonstrate how it corresponds to a non-trivial topological phase characterized by half-integer topological invariants. Upon mapping this system onto an equivalent 2D lattice decorated with a central line of defects, we provide indirect evidence of a refined bulk-boundary correspondence by relating the half-integer bulk invariants with the presence of a single in-gap edge state in the spectrum of the mapped lattice.
The rest of the paper is organized as follows. In Sec. II, we provide a brief overview on the theory of non-orthogonal bases, with an emphasis on how to project an operator onto such bases, and how to orthonormalize them by means of the Löwdin transformation [29]. In Sec. III, we introduce the model considered in this study, namely the diamond chain threaded by a finite flux per plaquette, in its clean version, deriving its solutions and the form of the overlap matrix for the CLSs spanning the zero-energy FB. In Sec. IV, we perturb the model by including weak local impurities in the top and bottom sites of the midchain plaquette, deriving the general analytical expression for the impurity states that come off the FB, using the techniques described in Sec. II. After that, we analyze in greater detail the particular cases of equal impurities, symmetric impurities, and single-impurity. For the latter, we also study its peculiar topological features, whose physical manifestation requires the introduction of an equivalent 2D lattice for its observation. Lastly, we present our conclusions in Sec. V.
II Basics of non-orthogonal bases
@@
II.1 Projection of operators onto a non-orthogonal subspace
@@ II.2 Symmetric orthonormalization
III Diamond chain with flux
@@
where $E_{\pm}(k)$ are dispersive for all $\phi$ except for $\phi=\pi$, where they become flat [see Fig. 1(b)]. An energy gap between the zero-energy FB $E_{0}$ and $E_{\pm}(k)$ opens at $k=\pi$ for $\phi\neq 0$, with a value of $E_{\text{gap}}=2t\sqrt{1+\cos\frac{\phi}{2}}$.
@@
Figure 1: (a) Diamond chain with $\phi$ flux per plaquette, distributed uniformly by the four hopping terms of each plaquette, with positive phase accumulated in the clockwise direction. Green shaded region represents the $q$th unit cell with sites $A_{q}$, $B_{q}$ and $C_{q}$, while blue and orange shaded regions represent the spatial width of adjacent CLSs |$\ket{\text{Loc},q+3}$ and |$\ket{\text{Loc},q+4}$, respectively, which overlap in the middle $B$ and $C$ sites. Onsite impurity potentials $\epsilon_{{}_{B}}$ and $\epsilon_{{}_{C}}$ are placed at the orange $B_{q+1}$ and light blue $C_{q+1}$ sites, respectively. (b) Bulk energy spectrum, in units of $t$, of the diamond chain for three representative values of the $\phi$ flux per plaquette. The zero-energy flat band is common to all $\phi$ values.
@@
yielding a tridiagonal real and symmetric matrix,
@@ IV Impurity states
Figure 2: Analytical energy curves (light blue and yellow curves) obtained from diagonalization of the effective two-state system in (35), as a function of the magnetic flux $\phi$ for (a) symmetric impurities $\epsilon_{{}_{B}}=-\epsilon_{{}_{C}}=\epsilon$, (b) equal impurities $\epsilon_{{}_{B}}=\epsilon_{{}_{C}}=\epsilon$, and (c) a single impurity, $\epsilon_{{}_{B}}=\epsilon$ and $\epsilon_{{}_{C}}=0$. Blue and red dots correspond to the numerically obtained energies of the two states of the FB that are affected by the impurity, for an open diamond chain with 53 plaquettes pierced by a $\phi$ flux, with the impurities placed at the $B$ and $C$ sites of the midchain plaquette in (a) and (b), and at its $B$ site in (c). Note that the overlap between analytical curves and numerical dots renders the former invisible for most of the spectrum in all three cases. (d)-(f) Probability amplitudes of the eigenstate marked by a blue circle in the corresponding cases above. Only the central part of the chain is shown. Golden bars (red dots) correspond to numerical (analytical) results. The left insets plot the same results in a semi-logarithmic scale and for a smaller portion of the chain. Sites without red dots are $A$ sites where the CLSs forming the analytical subspace have zero weight. The right insets depict the impurities at the midchain plaquette, with green (purple) color indicating an $\epsilon$ ($-\epsilon$) value for the impurity at that site. The blue circle (red square) in (c) also corresponds to the state shown on the top (bottom) of Fig. 5.
Since the two-state dual basis {|$\{\ket{\text{Loc},q^{*}},\ket{\text{Loc},q+1^{*}}\}$ is itself non-orthogonal, we first need to orhogonalize it through the Löwdin transformation [29, 35], as outlined in Sec. II.2. The first step requires us to find the form of $\mathbf{S}_{*}^{-\frac{1}{2}}$, where $\mathbf{S}_{*}$ is the overlap matrix written in this basis. Using (1), it is straightforward to show that the elements of $\mathbf{S}_{*}$ can be expressed as
@@
IV.1 Symmetric impurities ($\epsilon_{{}_{B}}=-\epsilon_{{}_{C}}=\epsilon$)
Figure 3: Snapshots of the amplitudes of the higher (lower) impurity state at the top (bottom) panel at different flux values, indicated on top, for a partial section of an open diamond chain with 53 complete plaquettes, with equal impurities $\epsilon=0.1$ in the middle one. The energy of both states at each flux value is given in Fig. 2(b).
For symmetric impurities, $\epsilon_{{}_{B}}=-\epsilon_{{}_{C}}=\epsilon$, where the impurity strength $\epsilon$ is given in units of the hopping parameter $t$, the effective impurity matrix in (35) reduces to
@@ IV.1 Symmetric impurities ($\epsilon_{{}_{B}}=-\epsilon_{{}_{C}}=\epsilon$)
In Fig. 2(a), we compare the analytical energies of (40) with numerical results as a function of the magnetic flux, for a chain with 53 plaquettes and symmetric local impurities at the top and bottom sites of the midchain plaquette [depicted in the right inset of Fig. 2(d)]. Two states with symmetric energies are lifted from the flat band for a finite magnetic flux per plaquette. Even though the perturbation is already quite large, $\epsilon=0.1$, that is, away for all $\phi$ from the $\epsilon\ll E_{\text{gap}}$ condition assumed in the projection of the impurity operator onto the flat band in (25), there is a nearly perfect match between analytical and numerical curves for the entire spectrum.
In Fig. 2(d), we plot the spatial profile of the eigenstate corresponding to the encircled positive energy at $\phi=\frac{\pi}{2}$ in Fig. 2(a). The analytical amplitude at each site, depicted as red dots, was computed by applying the general expression in (37), with the $\mathbf{U}$ matrix for the symmetric impurities case given in (41), and agrees very well with the numerical one, represented by the golden bar at the corresponding site. The inset shows the same results but in a semi-logarithmic scale. Sites without red dots correspond to A sites where the localized states have zero weight, according to (15)-(16). The negligible numerical weight at these sites comes from higher-order terms that couple the perturbed states with the top and bottom dispersive bands.
IV.2 Equal impurities ($\epsilon_{{}_{B}}=\epsilon_{{}_{C}}=\epsilon$)
@@
In Fig. 2(b), we compare the analytical energies of (43) with numerical results as a function of the magnetic flux for the impurity distribution depicted in the right inset of Fig. 2(e). Two positive energy states are lifted from the FB for a finite magnetic flux per plaquette, and become degenerate with $E=0.5\epsilon$ at $\phi=\pi$. Since $E_{\text{gap}}\to 0$ when $\phi\to 0\ (\text{mod}\ 2\pi)$, near the edges of the spectrum higher-order couplings to the dispersive bands can no longer be safely neglected, and the analytical treatment breaks down in that region. However, this region is still very small for $\epsilon=0.1$, and shrinks even further with decreasing $\epsilon$. The spatial amplitudes of the eigenstate with the energy marked by the blue circle in Fig. 2(b) are depicted in Fig. 2(e). As for the symmetric case above, a very good agreement is found between the numerical and analytical results.
IV.2.1 Disorder
@@
The localization length of the impurity states is controlled by the flux, as illustrated in Fig. 3. Having established numerically the validity of the analytical solutions considerably away from the $\epsilon\ll 1$ limit assumed in their derivation (the impurity strength is $\epsilon=0.1$ in Fig. 2, and even higher values still yield good agreement with numerical results), one can study the effect of small disorder on the impurity states. More specifically, we will consider diagonal disorder on an open diamond chain with 101 plaquettes threaded by a $\phi$ flux, by setting the onsite potential at each site to a random value taken from a uniform distribution in the interval $[-\frac{\delta}{2},\frac{\delta}{2}]$, with $\delta=0.02$ the disorder strength, with a standard deviation given by $\sigma_{\text{ud}}=\frac{\delta}{\sqrt{12}}\approx 5.8\times 10^{-3}$. For the sites at the midchain plaquette hosting the impurities, the disorder is added on top of $\epsilon=0.1$. We perform 20000 disorder realizations for each flux value and track the energies of the impurity states, from where we construct their respective histograms, as exemplified for the higher one at $\phi=\frac{9\pi}{20}$ in the inset of Fig. 4(b). Remarkably, the underlying uniform distribution of the disorder generates a normal distribution for the energies of the impurity states, whose mean approaches their disorder-free values given in Fig. 2(b) with increasing disorder realizations. This result is consistent with the central limit theorem (CLT), i.e., for each realization the impurity states effectively feel an averaged disorder related to their localization length, that is, to their spatial extension.
@@
Figure 4: (a) Standard deviation, as a function of the flux, of the energy of both impurity states for an open diamond chain with 101 plaquettes, with equal impurities $\epsilon=0.1$ in the middle one, taken from 20000 realizations of diagonal disorder at each point, with disorder strength $\delta=0.02$. (b) Effective number of sites involved in the disorder averaging effect, computed from the corresponding standard deviation values in (a) in accordance with the CLT. The inset shows the energy histogram for the encircled higher-energy state at $\phi=0.45\pi$, with the dashed orange curve corresponding to the Gaussian fit from which the standard deviation is extracted.
@@
In Fig. 4(a), we plot the standard deviation of the Gaussian fit of the energy distributions of both impurity states as a function of the flux . The flux was taken in $\frac{\pi}{20}$ increments and restricted to the $[\frac{3\pi}{20},\frac{15\pi}{20}]$ interval, since near $\phi=0,\pi$ the energy separation of the states is of the order of $\delta$ [see numerical results in Fig. 2(b) in these regions], which leads to a reordering of the states and to an overlap of their corresponding energy distributions, preventing their independent analysis. According to the CLT, the standard deviation of the normal distribution is related to the sample size, which in our case corresponds to the effective number of sites $N_{e}$ over which the disorder is averaged for each impurity state, through $\sigma=\frac{\sigma_{\text{ud}}}{\sqrt{N_{e}}}$, with $\sigma_{\text{ud}}$ given above. With this formula, and from the $\sigma$ values given in Fig. 4(a), we plot in Fig. 4(b) the flux dependence of $N_{e}$ for both impurity states. Interestingly, it can be seen that they follow opposite tendencies with increasing flux: while $N_{e}$ decreases exponentially for the lower state, it increases monotonously for the higher state.
@@
The evolution of the spatial profile of the impurity states with flux, shown in Fig. 3, can be used to explain their different response to disorder. The lower state becomes globally more and more localized until it becomes a CLS at $\phi=\pi$, which agrees with its decreasing $N_{e}$ in Fig. 4(b), since the disorder is being averaged over fewer sites, therefore generating a higher effective onsite disorder on the lower state. The higher state, on the other hand, has dominant contributions from the impurity sites at lower fluxes (see central peaks in the top panel of Fig. 3 with growing weight for decreasing flux). The weights at these sites are in an anti-symmetric configuration that becomes the full eigenstate (a CLS) at $\phi=0$, while they are in a symmetric configuration for the lower state. At zero flux, applying the same rotation to the $B$ and $C$ sites at each plaquette has been shown to decouple the anti-symmetric state, while the symmetric one connects with adjacent $A$ sites to form a uniform chain [40]. As flux is increased, the anti-symmetric configuration of impurity sites couples to the uniform chain with growing strength [41] and, as with the lower state, the higher state develops a spatially extended profile that becomes increasingly localized (see the evolution of the components at sites other than the central impurity ones in the top panel of Fig. 3). Conversely, the peaks at the impurity sites, which are the main contributors to the number of effective sites $N_{e}$ of the averaged disorder, continuously decrease with increasing flux, that is, spread their weight to the other sites. This effect, which tends to increase $N_{e}$, overpowers the localization effect of the other components, which tends to decrease $N_{e}$ (as for the lower state), leading to the net increase of $N_{e}$ with flux seen in Fig. 4(b).
@@
IV.3 Single impurity ($\epsilon_{{}_{B}}=\epsilon,\epsilon_{{}_{C}}=0$)
@@
where the formula for $\theta$ given in (21) was used in the last line. Both energy bands are flat and, in particular, we have $E_{-}=0$, as shown in Fig. 2(c), meaning that, when a single impurity is present, the system can always rotate the subspace of dual localized states in order to generate a state, namely |$\ket{\tilde{\varphi}_{-}}$ in (47), with zero weight at the impurity site, $\innerproduct{B_{q+1}}{\tilde{\varphi}_{-}}=0$, such that it remains degenerate with the other FB states. This is illustrated in Fig. 5, where a partial profile of the |$\ket{\tilde{\varphi}_{+}}$ and |$\ket{\tilde{\varphi}_{-}}$ eigenstates, computed through (47) for $\phi=\frac{\pi}{2}$, is depicted [the complete profile of |$\ket{\tilde{\varphi}_{+}}$ is shown in Fig. 2(f), which agrees very well with numerics]. It can be seen that the weight of the state at the impurity site is maximum for |$\ket{\tilde{\varphi}_{+}}$ and zero for |$\ket{\tilde{\varphi}_{-}}$.
Figure 5: Partial profile of the |$\ket{\tilde{\varphi}_{+}}$ and |$\ket{\tilde{\varphi}_{-}}$ eigenstates with energies marked with a blue circle and a red square in Fig. 2(c), respectively, and where the radius of each circle represents the amplitude of the wavefunction at the respective site and the color represents its phase, coded by the color bar at the right. The arrow indicates the $B_{q+1}$ site where the $\epsilon_{{}_{B}}=\epsilon=0.1$ impurity is placed, and where |$\ket{\tilde{\varphi}_{+}}$ (|$\ket{\tilde{\varphi}_{-}}$) has maximum (zero) amplitude.
IV.3.1 Topological characterization
The topological characterization of the single-impurity case reveals interesting new features. Let us first summarize the usual approach to the topological characterization of gapped 1D two-band models. The relevant topological invariant for characterizing the energy gap is the Zak phase of the lower band, which can be expressed as
@@ IV.3.1 Topological characterization
Figure 6: Phase $\vartheta$ of the lowest energy eigenstate as a function of the magnetic flux $\phi$ (momentum $k$) for the single-impurity model (SSH model). The color of each curve matches the color of the corresponding horizontal axis. For the SSH model, the ratio $t_{1}/t_{2}=2$ ($t_{1}/t_{2}=0.5$) was used for the topologically (non-)trivial case.

### v2

@@ Abstract
Flat band (FB) systems, featuring dispersionless energy bands, have garnered significant interest due to their compact localized states (CLSs). However, a detailed account on how local impurities affect the physical properties of overlapping CLSs is still missing. Here we study a diamond chain with a finite magnetic flux per plaquette that exhibits a gapped midspectrum FB with non-orthogonal CLSs, and develop a framework for projecting operators onto such non-orthogonal bases. This framework is applied to the case of an open diamond chain with small local impurities in the midchain plaquette, and analytical expressions are derived for FB states influenced by these impurities. For equal impurities in top and bottom sites under diagonal disorder, we show how the impurity states experience an averaged disorder dependent on their spatial extension, leading to enhanced robustness against disorder. For a single impurity, an exotic topological phase with a half-integer winding number is discovered, which is linked to a single in-gap edge state under open boundary conditions. Numerical simulations validate the analytical predictions.
Introduction
About a decade ago, a renewed interest in flat band (FB) systems [1] gave way to a new cycle of studies devoted to the topic (a more detailed chronology can be found in [2]), and it has no signs of slowing down. In this recent wave of studies, the focus has gradually shifted towards the compact localized states (CLSs) that span most types of FBs. For one-dimensional (1D) models, it has been shown that these CLSs form an orthogonal set when the FB projector is strictly local, and a non-orthogonal one otherwise [3, 4]. Based on the form of the CLSs, FB generators were constructed for 1D [5, 6] and higher-dimensional [7, 8, 9, 10, 11] models.
@@ Introduction
Here, we address the case of a diamond chain with a finite flux per plaquette, and introduce a perturbation in the form of weak local potentials in the midchain plaquette, as schematically illustrated in Fig. 1(a). By projecting this impurity perturbation onto the gapped zero-energy FB, only two partially overlapping CLSs have finite weight at the impurity sites. However, due to the non-orthogonal nature of adjacent CLSs, we show that the analytical treatment of the system requires one to reexpress the problem in terms of its dual basis states. This greatly simplifies the problem, by reducing it to the solutions of a decoupled dimer in a non-orthogonal dual subspace. We then derive the analytical solutions for the impurity states that are lifted from the FB as a function of the magnetic flux, focusing on the cases of two equal impurities, two opposite impurities, and a single-impurity, where near-perfect agreement with numerical results is observed. Finally, for the single-impurity case, we further demonstrate how it corresponds to a non-trivial topological phase characterized by half-integer topological invariants. Upon mapping this system onto an equivalent 2D lattice decorated with a central line of defects, we provide indirect evidence of a refined bulk-boundary correspondence by relating the half-integer bulk invariants with the presence of a single in-gap edge state in the spectrum of the mapped lattice.
Results
Basics of non-orthogonal bases
@@ Results
Projection of operators onto a non-orthogonal subspace
@@
Diamond chain with flux
@@
where $E_{\pm}(k)$ are dispersive for all $\phi$ except for $\phi=\pi$, where they become flat [see Fig. 1(b)]. An energy gap between the zero-energy FB $E_{0}$ and $E_{\pm}(k)$ opens at $k=\pi$ for $\phi\neq 0$, with a value of $E_{\text{gap}}=2t\sqrt{1-\cos\frac{\phi}{2}}$.
@@
Figure 1: Diamond chain. (a) Diamond chain with $\phi$ flux per plaquette, distributed uniformly by the four hopping terms of each plaquette, with positive phase accumulated in the clockwise direction. Green shaded region represents the $q$th unit cell with sites $A_{q}$, $B_{q}$ and $C_{q}$, while dashed blue and dotted orange regions represent the spatial width of adjacent CLSs |$\ket{\text{Loc},q+3}$ and |$\ket{\text{Loc},q+4}$, respectively, which overlap in the middle $B$ and $C$ sites. Onsite impurity potentials $\epsilon_{{}_{B}}$ and $\epsilon_{{}_{C}}$ are placed at the orange $B_{q+1}$ and light blue $C_{q+1}$ sites, respectively. (b) Bulk energy spectrum, in units of $t$, of the diamond chain for three representative values of the $\phi$ flux per plaquette. The zero-energy flat band is common to all $\phi$ values.
@@
yielding a tridiagonal real and symmetric Toeplitz matrix,
@@
Figure 2: Impurity states. Analytical energy curves (green and yellow curves) obtained from diagonalization of the effective two-state system in (35), as a function of the magnetic flux $\phi$ for (a) opposite impurities $\epsilon_{{}_{B}}=-\epsilon_{{}_{C}}=\epsilon$, (b) equal impurities $\epsilon_{{}_{B}}=\epsilon_{{}_{C}}=\epsilon$, and (c) a single impurity, $\epsilon_{{}_{B}}=\epsilon$ and $\epsilon_{{}_{C}}=0$, with $\epsilon=0.1t$. Blue and red dots correspond to the numerically obtained energies of the two states of the FB that are affected by the impurity, for an open diamond chain with 53 plaquettes pierced by a $\phi$ flux, with the impurities placed at the $B$ and $C$ sites of the midchain plaquette in (a) and (b), and at its $B$ site in (c). (d)-(f) Probability amplitudes of the eigenstate marked by a blue circle in the corresponding cases above. Only the central part of the chain is shown. Golden bars (red dots) correspond to numerical (analytical) results. The left insets plot the same results in a semi-logarithmic scale and for a smaller portion of the chain. Sites without red dots are $A$ sites where the CLSs forming the analytical subspace have zero weight. The right insets depict the impurities at the midchain plaquette, with green (purple) color indicating an $\epsilon$ ($-\epsilon$) value for the impurity at that site.
Since the two-state dual basis {|$\{\ket{\text{Loc},q^{*}},\ket{\text{Loc},q+1^{*}}\}$ is itself non-orthogonal, we first need to orhogonalize it through the Löwdin transformation [34, 35], as outlined in the “Symmetric orthonormalization” subsection of the Results. The first step requires us to find the form of $\mathbf{S}_{*}^{-\frac{1}{2}}$, where $\mathbf{S}_{*}$ is the overlap matrix written in this basis. Using (1), it is straightforward to show that the elements of $\mathbf{S}_{*}$ can be expressed as
@@
Opposite impurities
Figure 3: Evolution of impurity states with flux. Snapshots of the amplitudes of the (a) higher and (b) lower impurity state at different flux values, indicated on top, for a partial section of an open diamond chain with 53 complete plaquettes, with equal impurities $\epsilon=0.1t$ in the middle one. The energy of both states at each flux value is given in Fig. 2(b).
For opposite impurities, $\epsilon_{{}_{B}}=-\epsilon_{{}_{C}}=\epsilon$, where the impurity strength $\epsilon$ is given in units of the hopping parameter $t$, the effective impurity matrix in (35) reduces to
@@
In Fig. 2(a), we compare the analytical energies of (40) with numerical results as a function of the magnetic flux, for a chain with 53 plaquettes and opposite local impurities at the top and bottom sites of the midchain plaquette [depicted in the right inset of Fig. 2(d)]. Two states with opposite energies are lifted from the flat band for a finite magnetic flux per plaquette. Even though the perturbation is already quite large, $\epsilon=0.1t$, that is, away for all $\phi$ from the $\epsilon\ll E_{\text{gap}}$ condition assumed in the projection of the impurity operator onto the flat band in (25), there is a nearly perfect match between analytical and numerical curves for the entire spectrum.
In Fig. 2(d), we plot the spatial profile of the eigenstate corresponding to the encircled positive energy at $\phi=\frac{\pi}{2}$ in Fig. 2(a). The analytical amplitude at each site, depicted as red dots, was computed by applying the general expression in (37), with the $\mathbf{U}$ matrix for the opposite impurities case given in (41), and agrees very well with the numerical one, represented by the golden bar at the corresponding site. The inset shows the same results but in a semi-logarithmic scale. Sites without red dots correspond to A sites where the localized states have zero weight, according to (15)-(16). The negligible numerical weight at these sites, which is at least an order of magnitude smaller than the adjacent B and C sites with highest weight, comes from higher-order terms that couple the perturbed states with the top and bottom dispersive bands. Therefore, as the impurity value is decreased to approach the $\epsilon\ll E_{\text{gap}}$ limit, the weight at the A sublattice decreases even further, relative to the weight at the BC sublattice.
Equal impurities
@@
In Fig. 2(b), we compare the analytical energies of (43) with numerical results as a function of the magnetic flux for the impurity distribution depicted in the right inset of Fig. 2(e). Two positive energy states are lifted from the FB for a finite magnetic flux per plaquette, and become degenerate with $E=0.5\epsilon$ at $\phi=\pi$. Since $E_{\text{gap}}\to 0$ when $\phi\to 0\ (\text{mod}\ 2\pi)$, near the edges of the spectrum higher-order couplings to the dispersive bands can no longer be safely neglected, and the analytical treatment breaks down in that region. However, this region is still very small for $\epsilon=0.1t$, and shrinks even further with decreasing $\epsilon$. The spatial amplitudes of the eigenstate with the energy marked by the blue circle in Fig. 2(b) are depicted in Fig. 2(e). As for the case of opposite impurities above, a very good agreement is found between the numerical and analytical results.
@@
Having established numerically the validity of the analytical solutions considerably away from the $\epsilon\ll E_{\text{gap}}$ limit assumed in their derivation (the impurity strength is $\epsilon=0.1t$ in Fig. 2, and even higher values still yield good agreement with numerical results), one can study the effect of small disorder on the impurity states. More specifically, we will consider diagonal disorder on an open diamond chain with 101 plaquettes threaded by a $\phi$ flux, by setting the onsite potential at each site to a random value taken from a uniform distribution in the interval $[-\frac{\delta}{2},\frac{\delta}{2}]$, with $\delta=0.02t$ the disorder strength, with a standard deviation given by $\sigma_{\text{ud}}=\frac{\delta}{\sqrt{12}}\approx 5.8\times 10^{-3}$. For the sites at the midchain plaquette hosting the impurities, the disorder is added on top of $\epsilon=0.1t$. We perform 20000 disorder realizations for each flux value and track the energies of the impurity states, from where we construct their respective histograms, as exemplified for the higher one at $\phi=\frac{9\pi}{20}$ in the inset of Fig. 4(b). The underlying uniform distribution of the disorder generates a normal distribution for the energies of the impurity states, whose mean approaches their disorder-free values given in Fig. 2(b) with increasing disorder realizations. This result is consistent with the central limit theorem (CLT), i.e., for each realization the impurity states effectively feel an averaged disorder related to their localization length, that is, to their spatial extension, which is controlled by the flux, as illustrated in Fig. 3.
@@
Figure 4: Disorder effects in the impurity subspace. (a) Standard deviation, as a function of the flux, of the energy of both impurity states for an open diamond chain with 101 plaquettes, with equal impurities $\epsilon=0.1t$ in the middle one, taken from 20000 realizations of diagonal disorder at each point, with disorder strength $\delta=0.02t$. (b) Effective number of sites involved in the disorder averaging effect, computed from the corresponding standard deviation values in (a) in accordance with the CLT. The inset shows the energy histogram for the encircled higher-energy state at $\phi=0.45\pi$, with the dashed orange curve corresponding to the Gaussian fit from which the standard deviation is extracted.
@@
In Fig. 4(a), we plot the standard deviation of the Gaussian fit of the energy distributions of both impurity states as a function of the flux . The flux was taken in $0.05\pi$ increments and restricted to the $[0.15\pi,0.75\pi]$ interval, since near $\phi=0,\pi$ the energy separation of the states is of the order of $\delta$ [see numerical results in Fig. 2(b) in these regions], which leads to a reordering of the states and to an overlap of their corresponding energy distributions, preventing their independent analysis. According to the CLT, the standard deviation of the normal distribution is related to the sample size, which in our case corresponds to the effective number of sites $N_{e}$ over which the disorder is averaged for each impurity state, through $\sigma=\frac{\sigma_{\text{ud}}}{\sqrt{N_{e}}}$, with $\sigma_{\text{ud}}$ given above. With this formula, and from the $\sigma$ values given in Fig. 4(a), we plot in Fig. 4(b) the flux dependence of $N_{e}$ for both impurity states. It can be seen that they follow opposite tendencies with increasing flux: while $N_{e}$ decreases exponentially for the lower state, it increases monotonously for the higher state.
@@
The evolution of the spatial profile of the impurity states with flux, shown in Fig. 3, can be used to explain their different response to disorder. The lower state becomes globally more and more localized until it becomes a CLS at $\phi=\pi$, which agrees with its decreasing $N_{e}$ in Fig. 4(b), since the disorder is being averaged over fewer sites, therefore generating a higher effective onsite disorder on the lower state. The higher state, on the other hand, has dominant contributions from the impurity sites at lower fluxes (see central peaks in the top panel of Fig. 3 with growing weight for decreasing flux). The weights at these sites are in an anti-symmetric configuration that becomes the full eigenstate (a CLS) at $\phi=0$, while they are in a symmetric configuration for the lower state. At zero flux, applying the same rotation to the $B$ and $C$ sites at each plaquette has been shown to decouple the anti-symmetric state, while the symmetric one connects with adjacent $A$ sites to form a uniform chain [40, 41]. As flux is increased, the anti-symmetric configuration of impurity sites couples to the uniform chain with growing strength [42] and, as with the lower state, the higher state develops a spatially extended profile that becomes increasingly localized (see the evolution of the components at sites other than the central impurity ones in the top panel of Fig. 3). Conversely, the peaks at the impurity sites, which are the main contributors to the number of effective sites $N_{e}$ of the averaged disorder, continuously decrease with increasing flux, that is, spread their weight to the other sites. This effect, which tends to increase $N_{e}$, overpowers the localization effect of the other components, which tends to decrease $N_{e}$ (as for the lower state), leading to the net increase of $N_{e}$ with flux seen in Fig. 4(b).
@@
As for the finite energy impurity states studied in the “Opposite impurities” subsection above and in the “Single impurity” subsection below, we have checked numerically that this averaging effect is less pronounced. Under the same kind of diagonal disorder and for different flux values, the energy histogram of these other impurity states yields in general a cross between a Gaussian and a uniform distribution, signaling a number of effective sites involved in the averaging effect close to one, such that the CLT cannot be used to determine this number. This can be explained from the dominant weight that these states always have at one of the impurity sites [see the central peak in Figs. 2(d) and (f)], which is where the perturbative effects of disorder are felt the most by these states. Therefore, the CLT only provides an accurate description of the averaging effect of diagonal disorder for the equal impurities case.
Single impurity
@@
where the formula for $\theta$ given in (21) was used in the last line. Both energy bands are flat (in $\phi$) and, in particular, we have $E_{-}=0$, as shown in Fig. 2(c), meaning that, when a single impurity is present, the system can always rotate the subspace of dual localized states in order to generate a state, namely |$\ket{\tilde{\varphi}_{-}}$ in (47), with zero weight at the impurity site, $\innerproduct{B_{q+1}}{\tilde{\varphi}_{-}}=0$, such that it remains degenerate with the other FB states. This is illustrated in Fig. 5, where a partial profile of the |$\ket{\tilde{\varphi}_{+}}$ and |$\ket{\tilde{\varphi}_{-}}$ eigenstates, computed through (47) for $\phi=\frac{\pi}{2}$, is depicted [the complete profile of |$\ket{\tilde{\varphi}_{+}}$ is shown in Fig. 2(f), which agrees very well with numerics]. It can be seen that the weight of the state at the impurity site is maximum for |$\ket{\tilde{\varphi}_{+}}$ and zero for |$\ket{\tilde{\varphi}_{-}}$.
Figure 5: Midchain profile of impurity states. Partial profile of the |$\ket{\tilde{\varphi}_{+}}$ and |$\ket{\tilde{\varphi}_{-}}$ eigenstates at $\phi=\frac{\pi}{2}$, where the radius of each circle represents the amplitude of the wavefunction at the respective site and the color represents its phase, coded by the color bar at the right. The arrow indicates the $B_{q+1}$ site where the $\epsilon_{{}_{B}}=\epsilon=0.1t$ impurity is placed, and where |$\ket{\tilde{\varphi}_{+}}$ (|$\ket{\tilde{\varphi}_{-}}$) has maximum (zero) amplitude.
We turn now to the topological characterization of the single-impurity case. But first let us summarize the usual approach to the topological characterization of gapped 1D two-band models. The relevant topological invariant for characterizing the energy gap is the Zak phase of the lower band, which can be expressed as
@@
Figure 6: Phase of impurity state. Phase $\vartheta$ of the lowest energy eigenstate as a function of the magnetic flux $\phi$ (momentum $k$) for the single-impurity model (SSH model). The color of each curve matches the color of the corresponding horizontal axis. For the SSH model, the ratio $t_{1}/t_{2}=2$ ($t_{1}/t_{2}=0.5$) was used for the topologically (non-)trivial case.